OpenStudy (anonymous):

A train leaves Little Rock and travels north at 85 kilometers per hour. Another train leaves one hour later and travels north at 95 kilometers per hour. How long after the first train leaves do the trains meet? How far did the trains travel?

OpenStudy (cwrw238):

the distance travelled by the 2 trains is the same ket first rain take t1 hours and second train t2 hours speed = distance / time 85 = d / t1 so d = 85t1 95 = d / t2 so d = 95t2 therefore 85t1 = 95t2 also t1 - t2 = 1 solve these 2 equations for t1, t2 and then find d

OpenStudy (anonymous):

Im not sure how to solve them @cwrw238

OpenStudy (cwrw238):

ok use substitution 85t1 = 95t2 so t1= 95t2 / 85 = 19t2 / 17 now plug in 19t2/17 for t2 in the equation t1-t2 = 1:- 19t2 / 17 - t2 = 1 2t2/17 = 1 2t2 = 17 t2 = 8.5 hours

OpenStudy (cwrw238):

and t1 = 1 + 8.5 = 9.5 hours so its 9.5 hours after the first train leaves when they meet and the ditance travelled = 85 * 9.5 = 807.5 miles